Find the amplitude, frequency and period of motion
Answer #1

1) amplitude = 0.25 m

frequency = (\pi/8)/2\pi = 1/16 Hz = 0.0625 Hz

Time period = 1/f = 16 s

2) position = 0.25cos(\pi/8) = 0.23097 m

velocity = -(\pi/32)sin(\pi/8) = -0.03757 m/s

accelaration = -0.23097 x (\pi/8)2 = -0.03562 m/s2